# How May U.S. Courts Scrutinize Their Recidivism Risk Assessment Tools? Contextualizing AI Fairness Criteria on a Judicial Scrutiny-based Framework

### v1

@@ Abstract
The AI/HCI and legal communities have developed largely independent conceptualizations of fairness. This conceptual difference hinders the potential incorporation of technical fairness criteria (e.g., procedural, group, and individual fairness) into sustainable policies and designs, particularly for high-stakes applications like recidivism risk assessment. To foster common ground, we conduct legal research to identify if and how technical AI conceptualizations of fairness surface in primary legal sources. We find that while major technical fairness criteria can be linked to constitutional mandates such as “Due Process” and “Equal Protection” thanks to judicial interpretation, several challenges arise when operationalizing them into concrete statutes/regulations. These policies often adopt procedural and group fairness but ignore the major technical criterion of individual fairness. Regarding procedural fairness, judicial “scrutiny” categories are relevant but may not fully capture how courts scrutinize the use of demographic features in potentially discriminatory government tools like RRA. Furthermore, some policies contradict each other on whether to apply procedural fairness to certain demographic features. Thus, we propose a new framework, integrating demographics-related legal scrutiny concepts and technical fairness criteria.
@@ Introduction
Recidivism risk assessment (RRA) tools, widely deployed models that predict an individual’s likelihood to re-offend following a criminal charge, raise ethical concerns surrounding fairness, transparency, and potential bias. These tools rely on quantifiable factors (e.g., prior criminal history, age at first offense) to produce a risk score that assists court officials in bail, sentencing, parole, and other correctional decisions.
@@
RQ1.a: Do technical AI fairness criteria for RRA have an enforceable legal basis under the U.S. Constitution and its interpretations?
RQ1.b: Are there any legal inconsistencies/challenges when such abstract fairness criteria are translated into concrete policies?
RQ2: How to develop a new legal framework to mitigate those inconsistencies/challenges, informed by established legal theories?
Two corresponding contributions follow. First, we analyzed primary sources of law, including the U.S. Constitution, case law interpreting the Constitution, as well as statutes and regulations at the state and federal levels. The goal is to explore a legal basis for technical fairness criteria, as well as challenges when operationalizing legally relevant concepts of fairness. Second, to address the challenges, we propose an integrated framework that combines technical fairness criteria and the judicial “scrutiny” concept (typically applied when the government treats people differently based on their demographics). This “scrutiny” concept is particularly suited to the RRA use case (compared to other high-stakes AI-assisted decision-making use cases such as loan approval (Goyal et al. 2024) or job application screening (Rigotti and Fosch-Villaronga 2024), which are typically undertaken by private-sector actors) because the U.S. judiciary has developed elaborate scrutiny standards to scrutinize discriminations by public-sector actors (e.g., state and federal government bodies) against private citizens. Our framework utilizes a common unit of analysis, scrutiny-worthiness, to help auditors decide on which fairness-related criteria and for which demographics they should prioritize their audit.
@@
For a high-level overview, Narayanan et al. (2024) conducted a comprehensive review of empirical research on the perceived fairness of AI-assisted decision-making in the organizational setting (e.g., employee performance review, hiring). Their work builds upon the theoretical framework of organizational justice by Colquitt (2001), which includes four dimensions: procedural fairness, interpersonal fairness, informational fairness (i.e., explainability), and distributive (i.e., outcome-based) fairness. However, only two of their four fairness dimensions (distributive fairness, and to a lesser extent, procedural fairness) are frequently discussed within the fairness literature of the AI community, as comprehensively reviewed by Pessach and Shmueli (2022). Furthermore, Ryan, Nadal, and Doherty (2023) interviewed roughly equal numbers of AI and HCI experts using procedural fairness and (four mathematical variations of) distributive fairness. They found an interesting theme among the HCI experts’ (but not the AI experts) responses: “Meeting a mathematical definition of fairness does not mean the model is ethical”. This finding suggests that, when humans interact with an AI prediction system to make decisions, i.e., HCI is involved, fairness research should extend beyond technical formulation to consider the social, cultural, and legal background where the system is deployed.
@@
Legal and technical AI/HCI research on conceptualizations of fairness have evolved mostly on their own, evidenced in the communities having studied similar or neighboring phenomena yet invented different terminologies and frameworks. For example, a well-received U.S. law review article on RRA fairness maps “disparate impact” to outcome fairness, and “disparate treatment” both to process (i.e. procedural) fairness and loosely to individual fairness (Mayson 2019). This mapping implies that individual fairness is not a subcategory of outcome fairness and that individual fairness is more related to procedural fairness than to group fairness. Both are inconsistent with the technical literature. Another seminal work is the Legal chapter (“Understanding United States anti-discrimination law”) in the Fair Machine Learning book by Barocas, Hardt, and Narayanan (2023). This book chapter, however, focuses its investigation of fairness concepts from the perspective of law review articles, and not primary sources of U.S. legal authority (e.g., case law). This is a methodology gap we seek to fill. While Yang and Dobbie (2020) combined thorough law review and statistical approaches, their work interprets the “Equal Protection” clause mostly with procedural fairness and may ignore outcome fairness criteria.
@@
As explained by Phillips (2019), in the U.S., there are four main sources of law at the federal level: the U.S. Constitution (the most binding authority), statutes (from the legislative branch), regulations (from the executive branch) which are more detailed implementation of statutes, and case law (from the judicial branch which follows a hierarchy in order of increasing bindingness: district courts, circuit courts of appeal, and the court of last resort or the Supreme Court) which interpret the Constitution and determine if any statutes or regulations are unconstitutional. Regarding case law, one important nuance is that a decision from a higher court is not binding for a lower court unless the lower court is within the (often geography-defined) jurisdiction of the higher court. For example, case law by the Second Circuit Court of Appeals are only binding for federal district courts within three states (Connecticut, New York, Vermont) and not for federal district courts located in other states. An analogous system of legal sources can be found at the state level in most U.S. states. The main principle of interaction between federal and state laws is that the rights that federal laws offer to individuals is a lower bound for state laws, i.e., states can grant their people more, but not fewer rights. Criminal law is a legal area reserved mostly for state laws, except for matters that cross the state borders, e.g., internet and airplane crimes, which fall under the federal jurisdiction (Samaha 2016). Therefore, in the context of RRA, there exist two parallel sets of state and federal laws for bail, sentencing, and parole decisions.
@@
For example, U.S. Supreme Court’s (SCOTUS) has recognized race and alienage as strict scrutiny features through several decisions: 1/ Korematsu v. United States, 323 U.S. 214 (1944): “all legal restrictions which curtail the civil rights of a single racial group are immediately suspect. […] courts must subject them to the most rigid scrutiny.” (U.S. Supreme Court 1944). 2/ Graham v. Richardson, 403 U.S. 365 (1971): “classifications based on alienage, like those based on nationality or race, are inherently suspect and subject to close judicial scrutiny” (U.S. Supreme Court 1971). 3/ Bernal v. Fainter, 467 U.S. 216 (1984): “a state law that discriminates on the basis of alienage can be sustained only if it can withstand strict judicial scrutiny” (U.S. Supreme Court 1984).
@@
The second difference is a “political function” exception, which enables state/federal government to justify alienage-based discrimination but not race-based discrimination. As summarized in Bernal v. Fainter, 467 U.S. 216 (1984), “We [SCOTUS] have, however, developed a narrow exception to the rule that discrimination based on alienage triggers strict scrutiny […] and applies to laws that exclude aliens from positions intimately related to the process of democratic self-government” (U.S. Supreme Court 1984). SCOTUS used this “political function” exception to avoid striking down state policies that exclude aliens (non-U.S. citizens) from public service jobs such as police officers in Foley v. Connelie (U.S. Supreme Court 1978), teachers in Ambach v. Norwick (U.S. Supreme Court 1979), and probation officers in Cabell v. Chavez-Salido (U.S. Supreme Court 1982).
@@
Suppose future empirical works, e.g., human studies with crowdworkers or court officials, can confirm that our proposed three scrutiny thresholds are consistent with humans’ fairness perception, our framework might help systematically determine whether procedural, group, and/or individual fairness criteria should be audited with respect to each demographic feature. For example, if both crowdworkers and experts come to a consensus that age should be used as an input feature for recidivism risk assessment (i.e. age is below or fails the procedural fairness threshold and thus the group fairness and the individual fairness thresholds of scrutiny too), our framework indicates that parity with respect to age groups does not matter, but age can be used to compute the similarity function for individual fairness.
@@ Discussion
In fact, judges have very little incentive to consider algorithmic recourse or avoid overreliance on RRA tools because, as long as a judge still makes the final bail/sentencing decision, it becomes a shield to protect their decision against accusations of fairness violation in future appeals. For example, in Brooks v. Commonwealth, although there is evidence that the trial judge favors the AI risk assessment-based sentencing recommendation and dismisses the shorter active sentence in the non-AI recommendation, the Court of Appeals of Virginia rules that ‘the trial court properly exercised its discretion’ without inquiring the level of reliance by the trial court on the AI risk assessment (Court of Appeals of Virginia 2004).
@@
Before one can find recourse for bad algorithmic decisions, one must be able to contest those. The contestability literature in HCI often assumes the one who should contest is the model user (Vaccaro et al. 2019). However, Lyons, Velloso, and Miller (2021) found that first, the primary goal of contestability is to protect individuals, and second, contestability in AI-assisted decision-making resembles contestability in human decision-making. Therefore, in RRA and other high-stakes contexts where the decision subjects (e.g., defendants) differ from the model users (e.g., judges), the primary focus of contestability design should be on the decision subjects because they have the most incentives to protect themselves and are separate from the human decision makers, i.e., tool users (so that research in AI-assisted decision-making contestability can draw analogies from human decision-making contestability).
@@
Rudin (2019) argues against using black-box RRA tools because even if one tries to interpret black-box models’ predictions with post hoc explanations (developing a second model to perturb the original model’s outcomes to explain it), those explanations are not necessarily faithful. Their advocating for inherently interpretable models in high-stakes contexts like RRA will enable easier audit of “procedural fairness” in our framework, as expert witnesses might investigate the source code to determine whether the model uses a demographic feature and with how much weight.
Both works by Lagioia, Rovatti, and Sartor (2023) and Chouldechova (2017) demonstrate the potential incompatibility of several group fairness metrics (“a system that is equally accurate for different groups may fail to comply with group-parity standards”; “disparate impact can arise when an RPI fails to satisfy the criterion of error rate balance”), informing our future work to explore other variations of our group fairness-based (yellow) scrutiny threshold and their relative orders in our framework.

### v2

@@ Abstract
The AI/HCI and legal communities have developed largely independent conceptualizations of fairness. This conceptual difference hinders the potential incorporation of technical fairness criteria (e.g., procedural, group, and individual fairness) into sustainable policies and designs, particularly for high-stakes applications like recidivism risk assessment. To foster common ground, we conduct legal research to identify if and how technical AI conceptualizations of fairness surface in primary legal sources. We find that while major technical fairness criteria can be linked to constitutional mandates such as “Due Process” and “Equal Protection” thanks to judicial interpretation, several challenges arise when operationalizing them into concrete statutes/regulations. These policies often adopt procedural and group fairness but ignore the major technical criterion of individual fairness. Regarding procedural fairness, judicial “scrutiny” categories are relevant but may not fully capture how courts scrutinize the use of demographic features in potentially discriminatory government tools like RRA. Furthermore, some policies contradict each other on whether to apply procedural fairness to certain demographic features. Thus, we propose a new framework, integrating U.S. demographics-related legal scrutiny concepts and technical fairness criteria, and contextualize it in three other major AI-adopting jurisdictions (EU, China, and India).
@@ Introduction
Recidivism risk assessment (RRA) tools, widely deployed models that predict an individual’s likelihood to re-offend following a criminal charge, raise ethical concerns surrounding fairness, transparency, and potential bias. These tools rely on quantifiable factors (e.g., criminal history, age at first offense) to produce a risk score that assists court officials in bail, sentencing, parole, and other correctional decisions.
@@
RQ1.a: Do technical AI fairness criteria for RRA have an enforceable legal basis under the U.S. Constitution?
RQ1.b: Are there any legal challenges when such abstract fairness criteria are translated into concrete policies?
RQ2: How to develop a new legal framework to mitigate those challenges, informed by established legal theories?
Two main contributions follow. First, we analyzed primary sources of law, including the U.S. Constitution, case law interpreting the Constitution, as well as statutes and regulations at the state and federal levels. The goal is to explore a legal basis for technical fairness criteria, as well as challenges when operationalizing legally relevant concepts of fairness. Second, to address the challenges, we propose an integrated framework that combines technical fairness criteria and the judicial “scrutiny” concept (typically applied when the government treats people differently based on their demographics). This “scrutiny” concept is particularly suited to the RRA use case (compared to other high-stakes AI-assisted decision-making use cases such as loan approval (Goyal et al. 2024) or job application screening (Rigotti and Fosch-Villaronga 2024), which are typically undertaken by private-sector actors) because the U.S. judiciary has developed elaborate scrutiny standards to scrutinize discriminations by public-sector actors (e.g., state and federal government bodies) against private citizens. Our framework utilizes a common unit of analysis, scrutiny-worthiness, to help auditors decide on which fairness-related criteria and for which demographics they should prioritize their audit.
@@
For a high-level overview, Narayanan et al. (2024) conducted a comprehensive review of empirical research on the perceived fairness of AI-assisted decision-making in the organizational setting (e.g., employee performance review, hiring). Their work builds upon the theoretical framework of organizational justice by Colquitt (2001), which includes four dimensions: procedural fairness, interpersonal fairness, informational fairness (i.e., explainability), and distributive (i.e., outcome-based) fairness. However, only two of their four fairness dimensions (distributive fairness, and to a lesser extent, procedural fairness) are frequently discussed within the fairness literature of the AI community, as comprehensively reviewed by Pessach and Shmueli (2022). Furthermore, Ryan, Nadal, and Doherty (2023) interviewed roughly equal numbers of AI and HCI experts using procedural fairness and (four mathematical variations of) distributive fairness. They found an interesting theme among the HCI experts’ (but not the AI experts) responses: “Meeting a mathematical definition of fairness does not mean the model is ethical”. This finding suggests that, when humans interact with an AI prediction system to make decisions, i.e., HCI is involved, fairness research should extend beyond technical formulation to consider the social and legal background where the system is deployed.
@@
Legal and technical AI/HCI research on conceptualizations of fairness have evolved mostly on their own, evidenced in the communities having studied similar or neighboring phenomena yet invented different terminologies and frameworks. For example, a well-received U.S. law review article on RRA fairness maps “disparate impact” to outcome fairness, and “disparate treatment” both to process (i.e., procedural) fairness and loosely to individual fairness (Mayson 2019). This mapping implies that individual fairness is not a subcategory of outcome fairness and that individual fairness is more related to procedural fairness than to group fairness. Both are inconsistent with the technical literature. Another seminal work is the Legal chapter (“Understanding United States anti-discrimination law”) in the Fair Machine Learning book by Barocas, Hardt, and Narayanan (2023). This book chapter, however, focuses its investigation of fairness concepts from the perspective of law review articles, and not primary sources of U.S. legal authority (e.g., case law). This is a methodology gap we seek to fill. While Yang and Dobbie (2020) combined thorough law review and statistical approaches, their work interprets the “Equal Protection” clause mostly with procedural fairness and may ignore outcome fairness criteria.
@@
As explained by Phillips (2019), in the U.S., there are four main sources of law at the federal level: the U.S. Constitution (the most binding authority), statutes (from the legislative branch), regulations (from the executive branch) which are more detailed implementation of statutes, and case law (from the judicial branch which follows a hierarchy in order of increasing bindingness: district courts, circuit courts of appeal, and the court of last resort or the Supreme Court) which interpret the Constitution and determine if any statutes or regulations are unconstitutional. Regarding case law, one important nuance is that a decision from a higher court is not binding for a lower court unless the lower court is within the (often geography-defined) jurisdiction of the higher court. For example, case law by the Second Circuit Court of Appeals are only binding for federal district courts within three states (Connecticut, New York, Vermont) and not for federal district courts located in other states. An analogous system of legal sources can be found at the state level in most U.S. states. The main principle of interaction between federal and state laws is that the rights that federal laws offer to individuals is a lower bound for state laws, i.e., states can grant their people more, but not fewer rights. Criminal law is a legal area reserved mostly for state laws, except for matters that cross the state borders, e.g., internet and airplane crimes, which fall under the federal jurisdiction (Samaha 2016). Therefore, in the context of RRA, there exist two parallel sets of state and federal laws.
@@
For example, U.S. Supreme Court’s (SCOTUS) decisions recognize race and alienage as strict scrutiny features: 1/ Korematsu v. United States, 323 U.S. 214 (1944): “all legal restrictions which curtail the civil rights of a single racial group are immediately suspect. […] courts must subject them to the most rigid scrutiny.” (U.S. Supreme Court 1944). 2/ Graham v. Richardson, 403 U.S. 365 (1971): “classifications based on alienage, like those based on nationality or race, are inherently suspect and subject to close judicial scrutiny” (U.S. Supreme Court 1971). 3/ Bernal v. Fainter, 467 U.S. 216 (1984): “a state law that discriminates on the basis of alienage can be sustained only if it can withstand strict judicial scrutiny” (U.S. Supreme Court 1984).
@@
The second difference is a “political function” exception, which enables state/federal government to justify alienage-based discrimination but not race-based discrimination. In Bernal v. Fainter, 467 U.S. 216 (1984), “We [SCOTUS] have, however, developed a narrow exception to the rule that discrimination based on alienage triggers strict scrutiny […] and applies to laws that exclude aliens from positions intimately related to the process of democratic self-government” (U.S. Supreme Court 1984). SCOTUS used this “political function” exception to avoid striking down state policies that exclude aliens (non-U.S. citizens) from public service jobs such as police officers in Foley v. Connelie (U.S. Supreme Court 1978), teachers in Ambach v. Norwick (U.S. Supreme Court 1979), and probation officers in Cabell v. Chavez-Salido (U.S. Supreme Court 1982).
@@
Figure 4: How the empirical findings by Nguyen et al. (2025) might be applied to our scrutiny-based framework.
@@
Suppose empirical works, e.g., human studies with laypeople or court officials, can confirm that our proposed three scrutiny thresholds are consistent with humans’ fairness perception, our framework might help systematically determine whether procedural, group, and/or individual fairness criteria should be audited with respect to each demographic feature.
One empirical work highly relevant to our framework is the human subjects experiment conducted by Nguyen et al. (2025), which tests whether including a demographic feature (race, sex/gender, or age) as available information for laypeople to evaluate pairwise similarity between criminal justice defendants would worsen those laypeople’s individual fairness judgement (i.e., making them more likely to judge opposite risk scores on similar defendants as fair). They conclude that the pairwise individual similarity function should include sex and age but ignore race. They also have a short survey question on procedural fairness and find that most laypeople prefer demographic features (race, sex/gender, and age) to be excluded from the RRA model inputs.
Figure 4 summarizes how the human subjects experiment by Nguyen et al. (2025) might further clarify our proposed framework. Regarding the “exclusion from model inputs” (yellow) threshold, most of their participants believe criminal history features should be included (i.e., below the yellow threshold) and a majority of participants think that age, sex/gender, and race should be excluded (i.e., above the yellow threshold). Regarding the “ignorance in individual similarity function” (red) threshold, their empirical results and analyses conclude that the pairwise individual similarity function can consider age and sex/gender (i.e., below the red threshold), but it should ignore race (i.e., above the red threshold). Note that due to the relative positions of the red and yellow thresholds compared to sex/gender and age (red >>> { sex/gender, age } >>> yellow), applying the transitive property of inequality, the red threshold is above the yellow threshold, partially validating our theoretically proposed ordering. Finally, the three (blue) scrutiny ranges from legal precedents (strict scrutiny, intermediate scrutiny, rational basis) are placed around each respective demographic feature. However, their results do not confirm whether the red threshold falls within the strict or intermediate scrutiny range.
As a caution, this is just one out of many possible empirical arrangements of our proposed framework. For example, another possible arrangement is that sex/gender and age be placed below the yellow (procedural fairness) threshold because Katsiyannis et al. (2018) found sex/gender and age to be predictive of recidivism. Therefore, we invite future empirical works to refine our judicial scrutiny-based framework of fairness criteria and demographic features.
International Perspectives: How our U.S. Law Findings Relate to Other Major Jurisdictions
We contextualize our U.S. legal research findings and proposed scrutiny-based framework in other jurisdictions where there is a strong AI market and evidenced/potential use of AI-assisted decision-making tools by government bodies.
The European Union
The EU AI Act categorizes RRA tools (”Individual criminal offence risk assessment or prediction”) as “unacceptable risk” and therefore bans it. However, according to Recital 42: “risk assessments carried out with regard to natural persons in order to assess the likelihood of their offending or to predict the occurrence of an actual or potential criminal offence based solely on profiling them or on assessing their personality traits and characteristics should be prohibited. In any case, that prohibition does not refer to or touch upon risk analytics that are not based on the profiling of individuals or on the personality traits and characteristics of individuals, such as AI systems using risk analytics to assess the likelihood of financial fraud by undertakings on the basis of suspicious transactions” (The European Parliament and Council 2024). This likely means that any AI-assisted decision making tools used by EU member states are prohibited from including any demographic features (e.g., race, sex, or age) as model input features, i.e., all demographic features are placed above the “exclusion from model inputs” threshold if we adapt our scrutiny-based framework to the EU context.
Prior to the EU AI Act, as reviewed by Weerts et al. (2023), four main directives lay out the theoretical framework for EU non-discrimination law, which also applies to algorithmic fairness considerations: the Race Equality Directive 2000/43/EC (Council of European Union 2000b), the Framework Equality Directive (Council of European Union 2000a), the Gender Equality Directives 2004/113/EC (Council of European Union 2004) and 2006/54/EC (Council of European Union 2006). This EU framework defines two types of fairness concerns: direct discrimination (”one person is treated less favourably than another is, has been or would be treated in a comparable situation on grounds of” a protected feature like race, sex/gender, and age) and indirect discrimination (“an apparently neutral provision, criterion or practice would put persons of a protected group at a particular disadvantage”). While the strictly forbidden “direct discrimination” directly reaffirms our procedural fairness observation from the EU AI Act that race, sex, or age should be excluded from model input features, the less restricted “indirect discrimination” under EU non-discrimination law hints at our “group parity required” scrutiny threshold. Interestingly, “indirect discrimination” might be legal if it “is objectively justified by a legitimate aim and the means of achieving that aim are appropriate and necessary”, which was interpreted as “if the same legitimate aim can be achieved through less discriminatory alternatives, those must be used” (Weerts et al. 2023; Tobler 2005). This “[no] less discriminatory alternatives” language of permissible “indirect discrimnation” under EU non-discrimination law is highly similar to the “least restrictive means” language in the U.S. “strict scrutiny” review standard against discriminatory government policies. If an AI model with disparate impact (e.g., violating demographic parity against certain races) can survive “indirect discrimination” challenges under EU law, it is likely to survive the “strict scrutiny” test under U.S. law too.
Regarding individual fairness, Weerts et al. (2023) find the “likes should be treated alike” Aristotelian formulation of justice to be the philosophical foundation of the EU “direct discrimination” doctrine. However, after thoroughly reviewing specific and operationalizable AI-related sources of EU law such as the General Data Protection Regulation, GDPR (The European Parliament and Council 2016), the Digital Services Act, DSA (The European Parliament and Council 2022), and a proposal version of the EU AI Act, Calvi and Kotzinos (2023) conclude that “the few references to fairness metrics existing in the EU legal framework hint at group fairness and not individual fairness.” This problem is similar to our finding that although SCOTUS linked the U.S. constitutional “Equal Protection” clause to the technical individual fairness concept, we find no concrete individual fairness standards in U.S. federal/state statutes or regulations.
China
Although Chinese courts do not have an RRA use case, some of their criminal courts use AI to estimate the “social harm” of the current case based on similar past cases from other defendants (instead of predicting future criminal cases from the same defendant). Judges may consider AI-estimated “social harm” as “a frame of reference for sentencing,” which is analogous to U.S. judges considering RRA risk scores in sentencing (Papagianneas and Junius 2023; Cui 2020).
Regarding technical/legal fairness criteria, unlike the procedural vs. outcome fairness dichotomy in American/European laws, Papagianneas and Junius (2023) synthesize a procedural vs. substantive fairness dichotomy after reviewing judicial documents published by China’s Supreme People’s Court (SPC) to guide lower courts. Substantive fairness in China has nothing to do with outcome-based (group/procedural) fairness, but its goals are 1/ sustaining legitimacy (of the ruling party), 2/ maintaining social stability, and 3/ user convenience. On the other hand, procedural fairness in China is often less about a set of legally allowed/forbidden features for model inputs, but more about the explainability angle (internal accountability, procedural consistency, and external visibility), e.g., whether AI-assisted decision makers may explain to decision subjects what features they consider and why. Furthermore, the expected response to potentially discriminatory actions by a government agency in China is often to report to that agency or a higher government body for a top-down solution, rather than through bottom-up litigation against the government, i.e., judicial scrutiny of government actions is rare (Cui, Cheng, and Wiesner 2019). Therefore, our judicial scrutiny-based fairness framework is irrelevant to China’s legal system.
India
The most important supplement to our framework in India’s context is to expand the set of demographic features for fairness consideration to include “caste”, a hereditary Hindu class feature which historically determines one’s access to financial, cultural, social, and many other types of capital in life (Sambasivan et al. 2021). If placed on our legal framework, caste and sex/gender are likely above the “group parity required” threshold due to “reservations”, a constitutionally mandated practice to reserve certain quotas to access public resources like education and bank loans for vulnerable groups such as historically marginalized castes and women (Basavaraju 2009; Sambasivan et al. 2021). This policy seeks to equalize access to resources across demographic groups and is therefore related to group fairness. A recent case decided by the Supreme Court of India, LT. Col Nitisha v. Union of India (2021), though not in the AI context, further reaffirms group fairness protection for women as the Court recognizes indirect discrimination by the government (e.g., disparate impact of a policy on women, even without proven discriminatory intent) to be unconstitutional, just like direct discrimination (Supreme Court of India 2021).
@@ Discussion
Judges have little incentive to consider algorithmic recourse or avoid overreliance on RRA tools because, as long as a judge still makes the final bail/sentencing decision, it becomes a shield to protect their decision against accusations of fairness violation in future appeals. In Brooks v. Commonwealth, although there is evidence that the trial judge favors the AI risk assessment-based sentencing recommendation and dismisses the shorter active sentence in the non-AI recommendation, the Court of Appeals of Virginia rules that ‘the trial court properly exercised its discretion’ without inquiring the level of reliance by the trial court on the AI risk assessment (Court of Appeals of Virginia 2004).
@@
Before one can find recourse for bad algorithmic decisions, one must be able to contest those. The contestability literature in HCI often assumes the one who should contest is the model user (Vaccaro et al. 2019). However, Lyons, Velloso, and Miller (2021) found that first, the primary goal of contestability is to protect individuals, and second, contestability in AI-assisted decision-making resembles contestability in human decision-making. Therefore, in RRA and other high-stakes contexts where the decision subjects (e.g., defendants) differ from the model users (e.g., judges), the primary focus of contestability design should be on the decision subjects because they have the most incentives to protect themselves and are separate from the decision makers, i.e., tool users (so that research on AI-assisted decision-making contestability can draw analogies from human decision-making contestability).
@@
Rudin (2019) argues against using black-box RRA tools because even if one tries to interpret black-box models’ predictions with post hoc explanations (developing a second model to perturb the original model’s outcomes to explain it), those explanations are not necessarily faithful. Their advocating for inherently interpretable models in high-stakes contexts like RRA will enable easier audit of “procedural fairness” in our framework, as expert witnesses might investigate the source code to determine whether the model uses a demographic feature and with how much weight. Both works by Lagioia, Rovatti, and Sartor (2023) and Chouldechova (2017) demonstrate the potential incompatibility of several group fairness metrics (“a system that is equally accurate for different groups may fail to comply with group-parity standards”; “disparate impact can arise when an RPI fails to satisfy the criterion of error rate balance”), informing our future work to explore other variations of our group fairness-based (yellow) scrutiny threshold and their relative orders in our framework.